A car covers first half of the distance between two places at a speed of 80km/hr and
the second half at 60km/hr. What is the average speed of the car?
1
Expert's answer
2020-06-08T10:25:21-0400

"v_{avr}=\frac{S/2+S/2}{t_1+t_2}=\frac{S/2+S/2}{S/(2v_1)+S/(2v_2)}=\frac{v_1v_2}{v_1+v_2}=\frac{80\cdot 60}{80+60}=34.3km/hr"